# A compatible finite element discretisation for moist shallow water equations

### v1

@@ 1 Introduction
There are a number of significant processes in a weather or climate model that are not captured by the fluid dynamical equations that describe the behaviour of the atmosphere. These unresolved physical processes are parametrised, and often just called ‘physics’, in contrast to the discretisation of the resolved flow which is known as the ‘dynamics’. The physics includes processes that occur on scales that are too small or fast to be resolved by the solution of the fluid equations, and non-fluid dynamical processes such as radiation. In a forecast model the physics must be incorporated with the fluid mechanics solution to allow these processes to influence the large-scale dynamics correctly. This is collectively known as physics-dynamics coupling, and, as described in, for example [1], the implementation of this coupling can have a significant impact on the model as well as presenting challenges for numerical methods. For example, physics parametrisations can generate small-scale or grid-scale variability (and/or short-time variability), and it is important for numerical methods to be able to handle this appropriately. A common feature of physics schemes is the presence of discontinuities or switch-like processes which are also challenging from a numerical point of view.
@@
The structure of the article is as follows: in Section 2 we give an overview of the three most commonly-used moist shallow water models and their use in the literature, which motivates the idea of using moist processes in shallow water as a way to advance understanding about coupling physics processes to simple dynamical models. Section 3 describes the general formulation for the moist shallow water models and how to recover the different formulations, including the new moist convective pseudo-thermal model, from the general one. This section thus provides us with the framework with which we can compare formulations to further understanding of the different modelling choices. Both the compatible finite element discretisation of the model and our time-stepping scheme are dealt with in Section 4, which provides details of the new discretisation and demonstrates how our approach to solving the equations is highly relevant in a contemporary numerical weather prediction paradigm. Finally, the test cases and results in Section 5 offer reference solutions and information on modelling choices, thus realising the aim of moist shallow water models for physics-dynamics coupling investigations by providing the understanding and the means to be able to use these models in moist physics problems.
@@ 2.1 The shallow water equations
$\frac{\partial{D}}{\partial{t}}+\nabla\cdot(\bm{u}D)=0,$ (2)
where $\bm{u}$ is horizontal velocity, $D$ is the thickness of the fluid layer, $g$ is acceleration due to gravity, $f$ is the Coriolis parameter and $B$ is the height of the bottom surface.
@@ 2.2 The moist convective shallow water equations
The earliest model in this style is the seminal work by Gill [3]. Here an equation for a new moisture variable is added to the shallow water equations and convection is modelled via a sink in the depth equation. The model is used in a dam break problem, a Rossby adjustment problem, and to model wave propagation through moist and dry regions. The Gill model has also been used in studies of tropical data assimilation in, for example, [34] and [35]. Similar models to the Gill framework were used in studies of precipitation fronts by Frierson et al. [36], Stechman and Majda [2], and Pauluis et al. [37]. These use a linear baroclinic rotating shallow water-type model with a vertically averaged moisture field, with convection parameterised by a non-linear precipitation term. This non-linear source term gives rise to precipitation front solutions. Fronts are studied in the limit of infinitely fast relaxation in [36] and shown to be realisable in finite relaxation time in [2]. Interactions between a precipitation front and an incident disturbance are studied in [37].
@@
Variations of these one- and two-layer moist convective models are used in a number of studies of tropical dynamics. These include investigations of the instabilities of tropical cyclone-like vortices ([11], [12], [15] and [18], for example) and studies of equatorial modons and the Madden-Julian oscillation ([14], [39], [17], and [20], [16] for example.)
The formulation of the moist convective model we use is that of Bouchut et al. [4]. Condensation is triggered when the moisture variable exceeds some saturation threshold and acts to lower the layer thickness, through a sink term proportional to the condensation in the depth equation. The model is known as moist convective because the moist response triggers a convective one:
@@
where $Q$ is the water vapour in the shallow layer (the specific humidity integrated over the layer). Though topography $B$ is not included in the original model we include it here. Moisture satisfies the relaxation relation of the type used in general circulation models [40]:
@@ 2.3 The moist convective thermal shallow water equations
We use the model in its single layer version in our formulation (written here with topography, $B$):
@@ 2.4 The moist thermal shallow water equations
The idea of modelling the heating effect of moisture in the shallow water system is also dealt with by Zerroukat and Allen in [7], though with a different approach. Rather than beginning with a version of the usual shallow water equations and adding moisture, they derive a moist shallow water system from the three-dimensional Boussinesq approximation of the hydrostatic Euler equations. Unlike the usual Boussinesq equations, density is allowed to vary slightly with temperature, meaning the addition to the standard Boussinesq equations of an advection equation for a temperature/density variable and the addition to the momentum equation of buoyancy-related terms. The moist shallow water equations are derived by integrating this moist Boussinesq set over the depth of a shallow fluid layer. Moisture exists in three forms; vapour, cloud water and rain, and a three-state phase change model provides two-way physics-dynamics feedback. The framework is used in this study to demonstrate various test problems in the style of the Williamson et al. test cases [25] (steady-state geostrophically balanced flow, flow over an isolated mountain, mid-latitude jets over two isolated mountains, and zonal flow over Earth orography). The model is also used by Santos and Peixoto in [44] in their study of the impact of local grid refinement on cloud and rain formation over the Andes mountains, and by Ferguson and Jabonowski in [45] to test the ability of adaptive mesh refinement to track and resolve sharp gradients and small-scale flow filaments. Written here with topography $B$, the model is given by:
@@
The new variable $\theta$ represents the depth-averaged deviation form a background temperature. Here $q^{(k)},k=1,2,...,M$ are the moisture fields (for example, water vapour, cloud, or rain), and $S_{q}^{(k)}$ are source terms due to phase changes between the fields. $S_{\theta}$ is a source term for temperature due to latent heat release from moisture phase changes. Phase changes are described using a three-state moist physics scheme. Whenever the water vapour exceeds a threshold dictated by a prescribed saturation function (here a function of $D$ and $\theta$), a fraction of the excess is converted to cloud, with a corresponding latent heat release that heats the atmosphere locally. When the atmosphere is sub-saturated and there is cloud present a fraction of the cloud evaporates with a local cooling effect. When cloud passes a certain threshold a fraction of the excess is converted to rain, again with a corresponding latent heat release.
@@
Framework constant $b$ $S_{D}$ $S_{b}$ physics
moist convective X one-way
moist convective thermal X one-way
moist thermal X X 3-state
Table 1: A summary of the variations between the three most commonly-used moist shallow water models in the literature. The moist convective model can be considered to have a constant thermal or buoyancy variable (denoted in the table by $b$). $S_{D}$ is the right-hand side or source term in the depth equation, and $S_{b}$ the same for the buoyancy or potential temperature equation. These source terms are either present (✓) or absent (X) in each equation set. Both the moist convective model and the moist convective thermal model are usually coupled to a two-state physics scheme where conversions between water vapour and cloud are irreversible, called one-way physics in the table. The moist thermal model is coupled to a three-state moist physics scheme with water vapour, cloud and rain, where phases changes between water vapour and cloud are reversible.
@@ 3 General moist shallow water formulation
The moist convective model, the moist convective thermal model and the moist thermal model can be viewed as, respectively, a model where moist thermodynamics feeds back on (1) the depth equation, (2) the depth equation and the velocity equation, and (3) the velocity equation only. With this paradigm in mind, we consolidate these three approaches into one flexible model, with source terms that vary depending on the choice of moisture feedback. This requires writing both models that include thermal effects with the same thermal variable. We take $b=g(1-\theta)$, which bears similarity to the relationship between the potential temperature and buoyancy in the Boussinesq equations in, for example, [28], [46], and [47] (noting that here $\theta$ is dimensionless, following Zerroukat and Allen [7]). We replace the temperature variable $\theta$ with $1-b/g$ everywhere in the moist thermal shallow water equations (Equations (11) - (14)), which, after some manipulation, become:
$\frac{\partial{D}}{\partial{t}}+\bm{\nabla}\cdot(\bm{u}D)=0,$ (16)
$\frac{\partial b}{\partial t}+(\bm{u}\cdot\bm{\nabla})b=-gS_{b},$ (17)
where $S_{b}$ is the source term for the buoyancy equation due to latent heat release from phase changes.
The advantage of defining this relationship between $\theta$ and $b$ is that the velocity equation in the moist thermal model now shares a right-hand side with the moist convective thermal model. This means that all three models are encapsulated by this equation set, where different choices for the source terms and the buoyancy recover each of the individual models. Thus the equation set that we will discretise in the following section is given by:
@@
where $B$ is the topography and $S_{D}$ is a source term for the depth equation that captures the convective effect of phase changes. The source terms $S_{D}$, $S_{b}$ and $S_{q}$ and the relationship between them come from the moist physics scheme. We have coupled the equations to two different moist physics schemes which are described below. The moisture variables $q^{(k)}$ represent mixing ratios, where $Q$ is the integral of $q$ over the depth of the shallow water layer, as in Zerroukat and Allen’s moist thermal shallow water model [7]. The mixing ratio variables obey the advective form of the transport equation, while the integrated variables Q would obey the conservative form of the transport equation. Moisture conservation could still be obtained using integrated variables through the inclusion of the depth field in the moisture transport equations, akin to the approach of [45] in their moist shallow water formulation and the approach of [48] in the compressible Euler equations.
@@
$q_{v}\rightarrow q_{v}-C,$ (25)
$q_{c}\rightarrow q_{c}+C,$ (26)
$D\rightarrow D-\beta_{1}C,$ (27)
$b\rightarrow b-\beta_{2}C,$ (28)
@@
$q_{v}<q_{sat}:C$ $=\text{min}\left[q_{c},\gamma_{v}\frac{q_{sat}-q_{v}}{\tau_{v}}\right],$ (29)
$q_{v}\rightarrow q_{v}+C,$ (30)
$q_{c}\rightarrow q_{c}-C,$ (31)
$D\rightarrow D+\beta_{1}C,$ (32)
$b\rightarrow b+\beta_{2}C,$ (33)
@@
$q_{c}>q_{precip}:C$ $=\gamma_{r}\frac{q_{c}-q_{precip}}{\tau_{r}},$ (34)
$q_{r}\rightarrow q_{r}+C,$ (35)
$q_{c}\rightarrow q_{c}-C,$ (36)
@@
$\begin{split}\Delta q_{v}&=\text{max}\left[0,\gamma_{v}\left(\frac{q_{v}-q_{sat}}{\tau_{v}}\right)\right],\\ \Delta q_{c}&=\text{min}\left[q_{c},\text{max}\left[0,\gamma_{v}\left(\frac{q_{sat}-q_{v}}{\tau_{v}}\right)\right]\right],\\ \Delta q_{r}&=\text{max}\left[0,\gamma_{r}\left(\frac{q_{c}-q_{precip}}{\tau_{v}}\right)\right],\\ S_{v}&=\Delta q_{c}-\Delta q_{v},\\ S_{c}&=\Delta q_{v}-\Delta q_{c}-\Delta q_{r},\\ S_{r}&=\Delta q_{r},\\ S_{D}&=\Delta q_{c}-\Delta q_{v},\\ S_{b}&=\Delta q_{c}-\Delta q_{v},\end{split}$ (37)
@@
Following Appendix A in Zerroukat and Allen [7], for the moist thermal frameworks (moist thermal, moist convective thermal and moist convective pseudo-thermal) the saturation function is dependent on both depth and buoyancy [^1] We introduce a factor of $gH$ to the saturation function given in Zerroukat and Allen [7] to give a dimensionless number.:
@@
Then substituting (43a) and (43b) into (42) gives:
$q_{sat}(D^{n}-\beta_{1}\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right],b^{n}-\beta_{2}\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right])=q^{n}_{v}-\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right].$ (44)
The goal is to re-arrange this to find $\gamma_{v}$. This can be achieved by assuming that
@@
With these assumptions, we have that
$\begin{split}q_{sat}(D^{n}-\beta_{2}\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right],b^{n}-\beta_{2}\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right])\approx q_{sat}(D^{n},b^{n})\\ -\frac{\partial q_{sat}}{\partial D}(D^{n},b^{n})\beta_{1}\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right]-\frac{\partial q_{sat}}{\partial b}(D^{n},b^{n})\beta_{2}\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right],\end{split}$ (46)
so that
$\begin{split}q_{sat}(D^{n},b^{n})-\frac{\partial q_{sat}}{\partial D}(D^{n},b^{n})\beta_{1}\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right]-\frac{\partial q_{sat}}{\partial b}(D^{n},b^{n})\beta_{2}\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right]\\ =q^{n}_{v}-\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right].\end{split}$ (47)
This can be re-arranged for $\gamma_{v}$ to give:
$\begin{split}\gamma_{v}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right]-\gamma_{v}\frac{\partial q_{sat}}{\partial D}(D^{n},b^{n})\beta_{1}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right]-\gamma_{v}\frac{\partial q_{sat}}{\partial b}(D^{n},b^{n})\beta_{2}\left[q_{v}^{n}-q_{sat}(D^{n},b^{n})\right]\\ =q_{v}^{n}-q_{sat}(D^{n},b^{n}),\end{split}$ (48)
which gives:
@@
$\frac{\partial q_{sat}}{\partial D}(D^{n},b^{n})=\frac{\partial}{\partial D}\left(\frac{q_{0}H}{D+B}e^{20(1-b/g)}\right)=-\frac{20}{g}q_{sat}(D^{n},b^{n}),$ (50)
@@
$\frac{\partial q_{sat}}{\partial b}(D^{n},b^{n})=\frac{\partial}{\partial b}\left(\frac{q_{0}H}{D+B}e^{20(1-b/g)}\right)=-\frac{1}{D+B}q_{sat}(D^{n},b^{n}).$ (51)
@@
Our choice for $\beta_{1}$ comes from Equation (8) in Vallis and Penn [20]. Taking $L=2.4\times 10^{6}\text{ J kg}^{-1}$, $c_{p}=1004\text{ J kg}^{-1}\text{K}^{-1}$, $Q_{a}=0.02$, $T_{0}=300\text{K}$ and $H=10000\text{m}$ gives $\beta_{1}$ of approximately 1600 m. $\beta_{2}$ is $gL$ where $L\approxeq 10$ is the pseudo-latent heat from Zerroukat and Allen [7]. Following Zerroukat and Allen [7], we take the threshold for rain $q_{precip}$ to be $1\times 10^{-4}$ and the conversion rate for rain $\gamma_{r}$ to be $1\times 10^{-3}$.

### v2

# Exploring forms of the moist shallow water equations using a new compatible finite element discretisation

@@ 1 Introduction
There are a number of significant processes in a weather or climate model that are not captured by the fluid dynamical equations that describe the behaviour of the atmosphere. These unresolved physical processes are parametrised, and often just called ‘physics’, in contrast to the discretisation of the resolved flow which is known as the ‘dynamics’. The physics includes processes that occur on scales that are too small or fast to be resolved by the solution of the fluid equations, and non-fluid dynamical processes such as radiation. In a forecast model the physics must be incorporated with the fluid mechanics solution to allow these processes to influence the large-scale dynamics correctly. This is collectively known as physics-dynamics coupling, and, as described in, for example [1], the implementation of this coupling can have a significant impact on the model as well as presenting challenges for numerical methods. For instance, physics parametrisations can generate small-scale or grid-scale variability (and/or short-time variability), and it is important for numerical methods to be able to handle this appropriately. A common feature of physics schemes is the presence of discontinuities or switch-like processes which are also challenging from a numerical point of view.
@@
The structure of the article is as follows: in Section 2 we give an overview of the three most commonly-used moist shallow water models and their use in the literature. Section 3 describes the general formulation for the moist shallow water models and how to recover the different formulations from the general one. This section thus provides us with the framework with which we can compare formulations. Both the compatible finite element discretisation of the model and our time-stepping scheme are dealt with in Section 4, which provides details of the new discretisation and demonstrates how our approach to solving the equations is highly relevant in a contemporary numerical weather prediction paradigm. Finally, the test cases and results in Section 5 offer representative solutions to compare with and information on modelling choices.
@@ 2.1 The shallow water equations
$\frac{\partial{D}}{\partial{t}}+\bm{\nabla}\cdot(\bm{u}D)=0,$ (2)
where $\bm{u}$ is horizontal velocity, $D$ is the thickness of the fluid layer, $g$ is acceleration due to gravity, $f$ is the Coriolis parameter, $\bm{\hat{k}}$ is the radial unit vector and $B$ is the height of the bottom surface.
@@ 2.2 The moist convective shallow water equations
The earliest model in this style is the seminal work by Gill [3]. Here an equation for a new moisture variable is added to the shallow water equations and convection is modelled via a sink in the depth equation. The model is used in a dam break problem, a Rossby adjustment problem, and to model wave propagation through moist and dry regions. The Gill model has also been used in studies of tropical data assimilation in, for example, [34] and [35]. Similar models to the Gill framework were used in studies of precipitation fronts by Frierson et al. [36], Stechman and Majda [2], and Pauluis et al. [37]. These use a linear rotating shallow water-type model with a vertically averaged moisture field, with convection parameterised by a non-linear precipitation term. This non-linear source term gives rise to precipitation front solutions. Fronts are studied in the limit of infinitely fast relaxation in [36] and shown to be realisable in finite relaxation time in [2]. Interactions between a precipitation front and an incident disturbance are studied in [37].
@@
Variations of these one- and two-layer moist convective models are used in a number of studies of tropical dynamics. These include investigations of the instabilities of tropical cyclone-like vortices ([11], [12], [15] and [18], for example) and studies of equatorial modons and the Madden-Julian oscillation ([14], [39], [17], and [20], [16], for example.)
The formulation of the moist convective model we use is that of Bouchut et al. [4]. Condensation is triggered when the moisture variable exceeds some saturation threshold and acts to lower the layer thickness, through a sink term proportional to the condensation in the depth equation. The model is known as moist convective because the moist response triggers a convective one, where $-\beta P$ represents a mass loss due to convection out of the shallow water layer:
@@
where $Q$ represents the amount of water vapour in the layer (the specific humidity integrated over the layer), and so obeys a conservative form of the transport equation. Though topography $B$ is not included in the original model we include it here. Moisture satisfies the relaxation relation of the type used in general circulation models [40]:
@@ 2.3 The moist convective thermal shallow water equations
We use the model in its single layer version in our formulation (written here with topography, $B$) where the variable $b$ represents a buoyancy field:
@@ 2.4 The moist thermal shallow water equations
The idea of modelling the heating effect of moisture in the shallow water system is also dealt with by Zerroukat and Allen in [7], though with a different approach. Rather than beginning with a version of the usual shallow water equations and adding moisture, they derive a moist shallow water system from the three-dimensional Boussinesq approximation of the hydrostatic Euler equations. Unlike the usual Boussinesq equations, density is allowed to vary slightly with temperature, meaning the addition to the standard Boussinesq equations of an advection equation for a temperature/density variable and the addition to the momentum equation of buoyancy-related terms. The moist shallow water equations are derived by integrating this moist Boussinesq set over the depth of a shallow fluid layer. Moisture exists in three forms; vapour, cloud water and rain, and a three-state phase change model provides two-way physics-dynamics feedback. The framework is used in this study to demonstrate various test problems in the style of the Williamson et al. test cases [25] (steady-state geostrophically balanced flow, flow over an isolated mountain, mid-latitude jets over two isolated mountains, and zonal flow over Earth orography). The model is also used by Santos and Peixoto in [44] in their study of the impact of local grid refinement on cloud and rain formation over the Andes mountains, and by Ferguson and Jablonowski in [45] to test the ability of adaptive mesh refinement to track and resolve sharp gradients and small-scale flow filaments. Written here with topography $B$, the model is given by:
@@
The new variable $\theta$ represents the depth-averaged deviation from a background temperature. Here $q^{(k)}$ where $k\in 1,2,3$ are the moisture fields (water vapour, cloud, and rain), and $S_{q}^{(k)}$ are source terms due to phase changes between the fields. In contrast to the variable $Q$ these moist fields are specific humidities, and so satisfy the advective form of the transport equation. $S_{\theta}$ is a source term for temperature due to latent heat release from moisture phase changes. Phase changes are described using a three-state moist physics scheme. Whenever the water vapour exceeds a threshold dictated by a prescribed saturation function (here a function of $D$ and $\theta$), a fraction of the excess is converted to cloud, with a corresponding latent heat release that heats the atmosphere locally. When the atmosphere is sub-saturated and there is cloud present a fraction of the cloud evaporates with a local cooling effect. When cloud passes a certain threshold a fraction of the excess is converted to rain, again with a corresponding latent heat release.
@@
Framework prognostic b $S_{D}$ $S_{b}$ physics
moist convective X X one-way
moist convective thermal one-way
moist thermal X 3-state
Table 1: A summary of the variations between the three most commonly-used moist shallow water models in the literature. The moist convective thermal model and the moist thermal model have prognostic thermal or buoyancy variables (denoted by $b$), while this variable can be considered to be constant in the moist convective model. $S_{D}$ is the right-hand side or source term in the depth equation, and $S_{b}$ the same for the buoyancy or potential temperature equation. These source terms are either present (✓) or absent (X) in each equation set. Both the moist convective model and the moist convective thermal model are usually coupled to a two-state physics scheme where conversions between water vapour and cloud are irreversible, called one-way physics in the table. The moist thermal model is coupled to a three-state moist physics scheme with water vapour, cloud and rain, where phases changes between water vapour and cloud are reversible.
@@ 3 General moist shallow water formulation
The moist convective model, the moist convective thermal model and the moist thermal model can be viewed as, respectively, a model where moist thermodynamics feeds back on (1) the depth equation, (2) the depth equation and the velocity equation, and (3) the velocity equation only. With this paradigm in mind, we consolidate these three approaches into one flexible model, with source terms that vary depending on the choice of moisture feedback. This requires writing both models that include thermal effects with the same thermal variable. We take $b=g(1-\theta)$, which bears similarity to the relationship between the potential temperature and buoyancy in the Boussinesq equations in, for example, [28], [46], and [47] (noting that here $\theta$ is dimensionless, following Zerroukat and Allen [7]), and we replace the temperature variable $\theta$ with $1-b/g$ everywhere in the moist thermal shallow water equations (Equations (11) - (14)).
The advantage of defining this relationship between $\theta$ and $b$ is that the velocity equation in the moist thermal model will share a right-hand side with the moist convective thermal model. This means that all three models are encapsulated by one equation set, where different choices for the source terms and the buoyancy recover each of the individual models. Thus the equation set that we will discretise in the following section is given by:
@@
where $B$ is the topography, $S_{D}$ is a source term for the depth equation that captures the convective effect of phase changes and $S_{b}$ is a source term for the buoyancy equation that captures the heating effect of phase changes. The source terms $S_{D}$, $S_{b}$ and $S_{q}$ and the relationship between them come from the moist physics scheme. We have coupled the equations to two different moist physics schemes which are described below. The moisture variables $q^{(k)}$ represent mixing ratios, where $Q$ is the integral of $q$ over the depth of the shallow water layer, as in Zerroukat and Allen’s moist thermal shallow water model [7]. The mixing ratio variables obey the advective form of the transport equation, while the integrated variables Q would obey the conservative form of the transport equation. Moisture conservation could still be obtained using integrated variables through the inclusion of the depth field in the moisture transport equations, akin to the approach of [45] in their moist shallow water formulation and the approach of [48] in the compressible Euler equations.
@@
$q_{v}\rightarrow q_{v}-\Delta tC,$ (21)
$q_{c}\rightarrow q_{c}+\Delta tC,$ (22)
$D\rightarrow D-\Delta t\beta_{1}C,$ (23)
$b\rightarrow b-\Delta t\beta_{2}C,$ (24)
@@
$q_{v}<q_{sat}:C$ $=\text{min}\left[\frac{q_{c}}{\Delta t},\gamma_{v}\frac{q_{sat}-q_{v}}{\tau_{v}}\right],$ (25)
$q_{v}\rightarrow q_{v}+\Delta tC,$ (26)
$q_{c}\rightarrow q_{c}-\Delta tC,$ (27)
$D\rightarrow D+\Delta t\beta_{1}C,$ (28)
$b\rightarrow b+\Delta t\beta_{2}C,$ (29)
@@
$q_{c}>q_{precip}:P$ $=\gamma_{r}\frac{q_{c}-q_{precip}}{\tau_{r}},$ (30)
$q_{r}\rightarrow q_{r}+\Delta tP,$ (31)
$q_{c}\rightarrow q_{c}-\Delta tP,$ (32)
@@
$\begin{split}\Delta q_{v}&=\text{max}\left[0,\gamma_{v}\left(\frac{q_{v}-q_{sat}}{\tau_{v}}\right)\right],\\ \Delta q_{c}&=\text{min}\left[\frac{q_{c}}{\Delta t},\text{max}\left[0,\gamma_{v}\left(\frac{q_{sat}-q_{v}}{\tau_{v}}\right)\right]\right],\\ \Delta q_{r}&=\text{max}\left[0,\gamma_{r}\left(\frac{q_{c}-q_{precip}}{\tau_{r}}\right)\right],\\ S_{v}&=\Delta q_{c}-\Delta q_{v},\\ S_{c}&=\Delta q_{v}-\Delta q_{c}-\Delta q_{r},\\ S_{r}&=\Delta q_{r},\\ S_{D}&=\Delta q_{c}-\Delta q_{v},\\ S_{b}&=\Delta q_{c}-\Delta q_{v}.\end{split}$ (33)
@@
Following Appendix A in Zerroukat and Allen [7], for the moist thermal frameworks (moist thermal, moist convective thermal and moist convective pseudo-thermal) the saturation function is dependent on both depth and buoyancy [^1] We introduce a factor of $gH$ to the saturation function given in Zerroukat and Allen [7] to give a dimensionless number, but note that we modify the scaling factor $q_{0}$ accordingly to give a saturation function the same size as that in [7].:
@@
We make these substitutions in (38) and then assume that:
@@
We expand the left-hand side of Equation (38) using a first-order Taylor expansion and rearrange to give:
@@
$\frac{\partial q_{sat}}{\partial D}(D^{n},b^{n})=\frac{\partial}{\partial D}\left(\frac{q_{0}H}{D+B}e^{20(1-b/g)}\right)=-\frac{1}{D+B}q_{sat}(D^{n},b^{n}),$ (42)
@@
$\frac{\partial q_{sat}}{\partial b}(D^{n},b^{n})=\frac{\partial}{\partial b}\left(\frac{q_{0}H}{D+B}e^{20(1-b/g)}\right)=-\frac{20}{g}q_{sat}(D^{n},b^{n}).$ (43)
@@
Our choice for $\beta_{1}$ comes from Equation (8) in Vallis and Penn [20]. Taking $L=2.4\times 10^{6}\text{ J kg}^{-1}$, $c_{p}=1004\text{ J kg}^{-1}\text{K}^{-1}$, $Q_{a}=0.02$, $T_{0}=300\text{K}$ and $H=10000\text{m}$ gives $\beta_{1}$ of approximately 1600 m. $\beta_{2}$ is $gL$ where $L\approxeq 10$ is the pseudo-latent heat from Zerroukat and Allen [7]. Following Zerroukat and Allen [7], we take the threshold for rain $q_{precip}$ to be $1\times 10^{-4}$ and the conversion rate for rain $\gamma_{r}$ to be $1\times 10^{-3}\text{ s}^{-1}$.